OpenStudy (anonymous):

Could someone show me the steps on how to solve cube 3 to the square root 27/y^6

OpenStudy (zepp):

\[\LARGE \sqrt[3]{\frac{27}{y^6}}\]

OpenStudy (zepp):

Use the property of \[\LARGE \sqrt[3]\frac{a}{b}=\frac{\sqrt[3]a}{\sqrt[3]b}\]

OpenStudy (zepp):

Can you try?

OpenStudy (zepp):

@jolliffbrett Can you try to apply this property of your question?

OpenStudy (anonymous):

i came up with 3/y^2 is that correct

OpenStudy (zepp):

Correct! :D

OpenStudy (anonymous):

Thanks,how about the square root of 9x^6/64y^3